### v3

# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

Abstract
The most widely used generative models map noise and data distributions by matching flows or scores. However, they struggle to incorporate partial observations and additional priors—something energy-based models (EBMs) handle elegantly by simply adding corresponding scalar energy terms. We address this issue by proposing Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs. Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. Our method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold. Furthermore, we leverage the method’s flexibility to introduce an interaction energy that supports diverse mode exploration, which we demonstrate in a controlled protein-generation setting. Our approach focuses on learning a scalar potential energy—without time-conditioning, auxiliary generators, or additional networks—which marks a significant departure from recent EBM methods. We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling across diverse domains.
1 Introduction
[…]
Figure 1: Trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching. We highlight several individual trajectories in red to illustrate their distinct behaviors. Both Action Matching and Flow Matching learn time-dependent transports and are not trained for traversing the data manifold. Conversely, EBMs and Energy Matching are driven by time-independent fields that can be iterated indefinitely, allowing trajectories to navigate across modes. While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
Many generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) - or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021). In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not explicitly capture the unconditional data score and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these noised manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For example, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
EBMs (Hopfield, 1982; Hinton, 2002; LeCun et al., 2006) provide an alternative approach for approximating the data distribution by learning a scalar-valued function $E(x)$ that specifies an unnormalized density $p(x)\propto\exp\left(-E(x)\right)$. Rather than explicitly mapping noise samples onto the data manifold, EBMs assign low energies to regions of high data concentration and high energy elsewhere. This defines a Boltzmann distribution from which one can sample, for example, via Langevin sampling. In doing so, EBMs explicitly retain the likelihood information in $E(x)$. This likelihood information can then be used in conditional generation (e.g., to solve inverse problems), possibly together with additional priors simply by adding their energy terms (Du and Mordatch, 2019). Moreover, direct examination of local curvature on the data manifold—allows the computation of local intrinsic dimension (LID) (an important proxy for data complexity)—whereas diffusion models can only approximate such curvature in the proximity of noise samples.
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity. {mdframed}[hidealllines=true,backgroundcolor=blue!5]
Contributions.
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1.
When samples lie far from the data manifold, they are efficiently transported toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on both CIFAR-10 and ImageNet generation in terms of fidelity, and compares favorably to flow-matching and diffusion models—without auxiliary generators or time-dependent EBM ensembles.
Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated, with an application to conditional protein generation. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models. [^1] The code will be publicly released upon acceptance.
2 Energy matching
In this section, we show how a scalar potential $V(x)$ can simultaneously provide an optimal-transport-like flow from noise to data while also yielding a Boltzmann distribution that explicitly captures the unnormalized likelihood of the data.
The Jordan–Kinderlehrer–Otto (JKO) scheme.
The starting point of our approach is the JKO scheme (Jordan et al., 1998), which is the basis of the success of numerous recent generative models (Xu et al., 2023; Terpin et al., 2024; Choi et al., 2024). The JKO scheme describes the discrete-time evolution of a probability distribution $\rho_{t}$ along energy-minimizing trajectories in the Wasserstein space,
$\rho_{t+\Delta t}=\arg\min_{\rho}\underbrace{\frac{W_{2}^{2}\left(\rho,\rho_{t}\right)}{2\Delta t}}_{\text{Transport Cost}}+\underbrace{\int V_{\theta}(x)\mathrm{d}\rho(x)}_{\text{Potential Energy}}+\underbrace{\varepsilon(t)\int\rho(x)\log\rho(x)\mathrm{d}x}_{\text{Internal Energy (-\text{Entropy)}}}.$ (1)
Here, $\theta$ denotes the learnable parameters of the scalar potential $V_{\theta}(x)$, and $\varepsilon(t)$ is a temperature-like parameter tuning the entropic term. The transport cost is given by the Wasserstein distance,
$W_{2}^{2}(\rho,\rho_{t})=\min_{\gamma\in\Gamma(\rho,\rho_{t})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\norm{x-x_{t}}^{2}\mathrm{d}\gamma(x,x_{t}),$ (2)
where $\Gamma(\rho,\rho_{t})$ is the set of couplings between $\rho$ and $\rho_{t}$, i.e., the set of probability distributions on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\rho$ and $\rho_{t}$. Here, $d$ is the dimensionality of the data. Henceforth, we call OT coupling any $\gamma_{t}$ that yields the minimum in (2). When $\gamma_{t}=(\mathrm{id},T)_{\#}\rho$, i.e., it is the pushforward of the map $x\mapsto(x,T(x))$ for some function $T$, we say that $T$ is an OT map from $\rho$ to $\rho_{t}$.
Differently from most literature, we consider $\varepsilon(t)$ to be dependent on time and study the behavior of Equation 1 as $t\to\infty$. To fix the ideas, consider, for instance, a linear scheduling:
$\varepsilon(t)=\begin{cases}0,&0\leq t<\operatorname{\tau^{\ast}},\\ \varepsilon_{\max}\frac{t-\operatorname{\tau^{\ast}}}{1-\operatorname{\tau^{\ast}}},&\operatorname{\tau^{\ast}}\leq t\leq 1,\\ \varepsilon_{\max},&t\geq 1.\end{cases}$ (3)
First-order optimality conditions.
We follow the approach in Terpin et al. (2024) and study (1) at each time $t$ via its first-order optimality conditions (Lanzetti et al., 2024, 2025):
$\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)+\varepsilon(t)\nabla_{x}\log\rho_{t}(x)=0,\qquad(x,y)\in\mathrm{supp}(\gamma_{t})$ (4)
where $\gamma_{t}$ is an OT plan between the distributions $\rho_{t}$ and $\rho_{t+\Delta t}$ and $\mathrm{supp}(\gamma_{t})$ is the support of $\gamma_{t}$. That is, this condition has to hold for all pairs of points in the support of $\rho_{t}$ and $\rho_{t+\Delta t}$ that are coupled by OT. Intuitively, analyzing (4) provides us with two key insights:
  1. 1.
    For times $t<\tau^{*}$, $\varepsilon(t)=0$ and (4) becomes
    $\frac{1}{\Delta t}(x-y)+\nabla_{x}V_{\theta}(x)=0\qquad(x,y)\in\mathrm{supp}(\gamma_{t}).$ (5)
    That is, the system is in an OT, flow-like, regime.
  2. 2.
    Near the data manifold, which we aim at modeling with the equilibrium distribution $\rho_{\text{eq}}$ of (1), $\rho_{t}\approx\rho_{\text{eq}}$ and, thus, for $t\gg 1$, $x\approx y$ for all $(x,y)\in\mathrm{supp}(\gamma_{t})$. Then, we can simplify (4) as
    $\varepsilon_{\max}\nabla_{x}\log\rho_{\mathrm{eq}}(x)=\nabla_{x}V_{\theta}(x)\quad\Longrightarrow\quad\rho_{\mathrm{eq}}(x)\propto\exp\left(-\tfrac{V_{\theta}(x)}{\varepsilon_{\max}}\right).$
    Thus, the equilibrium distribution is described by an EBM, $\exp\left(-E(x)\right)$, with $E(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$.
Our approach in a nutshell.
Combining the two insights above, we propose a generative framework that combines OT and EBMs to learn a time-independent scalar potential $V_{\theta}(x)$ whose Boltzmann distribution,
$\rho_{\mathrm{eq}}(x)\propto\exp\left(-\frac{V_{\theta}(x)}{\varepsilon_{\max}}\right),$ (6)
matches $\rho_{\mathrm{data}}$. To transport samples efficiently from noise $\rho_{0}$ to $\rho_{\mathrm{eq}}\approx\rho_{\mathrm{data}}$, we use two regimes:
  • Away from the data manifold: $\varepsilon\approx 0$. The flow is deterministic and OT-like, allowing rapid movement across large distances in sample space.
  • Near the data manifold: $\varepsilon\approx\varepsilon_{\max}$. Samples diffuse into a stable Boltzmann distribution, properly covering all data modes.
By combining the long-range transport capability of flows with the local density modeling flexibility of EBMs, we achieve tractable sampling and explicit likelihood of samples $V_{\theta}(x)$; see Figure 1.
2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section 2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium (see Section 2.1.2).
2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
We begin by constructing a global velocity field $\nabla_{x}V_{\theta}(x)$ that carries noise samples $\{x_{0}\}$ to data samples $\{x_{\mathrm{data}}\}$ with minimal detours. For this, we consider geodesics in the Wasserstein space (Ambrosio et al., 2008). Practically, we compute the OT coupling $\gamma^{*}$ between two uniform empirical probability distributions, one supported on a mini-batch $\{x_{\text{data}}^{(j)}\}_{j=1}^{B}$ of the data, and one supported on a set of noise samples with the same cardinality $B$. These samples are drawn from an easy-to-sample distribution; in our case, a Gaussian. Since the probability distributions are uniform and empirical with the same number of samples, a transport map $T$ is guaranteed to exist (Ambrosio et al., 2008).
Remark 2.1 (OT solver).
Depending on the approach used to compute the OT coupling, one may find an OT map or not. Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map. In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs $\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)$ with $\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}$. In our experiments, we used the POT solver (Flamary et al., 2021) and we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023).
Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation $x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)}$, which is a point along the geodesic (Ambrosio et al., 2008). The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$. For this, we define the loss:
$\operatorname{\mathcal{L}_{\mathrm{OT}}}=\mathbb{E}^{\begin{subarray}{c}x_{\mathrm{data}}\in\mathcal{D}\\ t\sim U(0,1)\end{subarray}}\left[\norm{\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T({x^{(i)}_{\mathrm{data}}})}^{2}\right].$
This objective can be interpreted as a flow-matching formulation under the assumption that the velocity field is both time-independent and given by the gradient of a scalar potential, thereby imposing a curl-free condition. This aligns naturally with OT, which also yields a velocity field of zero curl—any rotational component would add unnecessary distance to the flow and thus inflate the transport cost without benefit. Our experimental evidence adds to the recent study by (Sun et al., 2025), in which the authors observed that time-independent velocity fields can, under certain conditions, outperform time-dependent noise-conditioned fields in sample generation.
Algorithm 1 Phase 1 (warm-up).
1:Initialize model parameters $\theta$
2:for iteration $0,1,\ldots$ do
3:    Sample $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    Sample $\{x_{0}^{(i)}\}_{i=1}^{B}\sim\mathcal{N}(0,I)$triangleright Random Gaussian samples
5:    $T\leftarrow\mathrm{OTsolver}(\{x_{0}^{(i)}\},\{x_{\mathrm{data}}^{(j)}\})$triangleright Compute OT map
6:    Sample $t\sim U(0,1)$
7:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}(\theta)\leftarrow\sum_{i}\|\nabla_{x}V_{\theta}(x_{t}^{(i)})+x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})\|^{2}$triangleright Loss function
8:    $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{OT}}(\theta)$triangleright Gradient update with learning rate $\alpha$
9:end for
10:return $\theta$$\triangleright$ Trained $\theta$
2.1.2 Contrastive objective $\operatorname{\mathcal{L}_{\mathrm{CD}}}$
Near the data manifold, $V_{\theta}(x)$ is refined so that $\rho_{\mathrm{eq}}(x)\propto\exp\left(-V_{\theta}(x)/\varepsilon_{\max}\right)$ matches the data distribution. We adopt the contrastive divergence loss described in EBMs (Hinton, 2002),
$\operatorname{\mathcal{L}_{\mathrm{CD}}}=\mathbb{E}_{x\sim p_{\text{data}}}\left[\frac{V_{\theta}(x)}{\varepsilon_{\max}}\right]-\mathbb{E}_{\tilde{x}\sim\mathrm{sg}(p_{\mathrm{eq}})}\left[\frac{V_{\theta}(\tilde{x})}{\varepsilon_{\max}}\right],$
where $\tilde{x}$ are “negative” samples of the equilibrium distribution induced by $V_{\theta}$. We approximate these samples using an MCMC Langevin chain (Welling and Teh, 2011). We split the initialization for negative samples: half begin at real data, and half begin at the noise distribution. This way, $V_{\theta}(x)$ forms well-defined basins around high-density regions while also shaping regions away from the manifold, correcting the generation. The $\mathrm{sg}(\cdot)$ indicates a stop-gradient operator, which ensures gradients do not back-propagate through the sampling procedure.
2.1.3 Dual objective
To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3). We define a sampling time $\operatorname{\tau_{s}}$. In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2. The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$. We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$. The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2. Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives. Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyper-parameters for each dataset are defined in Appendix D.
Algorithm 2 Phase 2 (main training).
1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from the warmed-up model (Algorithm 1)
2:for iteration $0,1,\ldots$ do
3:    Sample mini-batch $\{x_{\mathrm{data}}^{(j)}\}_{j=1}^{B}\sim\mathcal{D}$triangleright Mini-batch of data samples
4:    $\operatorname{\mathcal{L}_{\mathrm{OT}}}\leftarrow$ Use lines 4–7 from Algorithm 1
5:    Initialize $\{x_{\mathrm{neg}}^{(0)}\}$ from noise and/or data $\triangleright$ Negative samples
6:    for $m=0,1,\ldots,M_{\text{Langevin}}$ do
7:         Sample $\eta\sim\mathcal{N}(0,I)$
8:         $x_{\mathrm{neg}}^{(m+1)}\leftarrow x_{\mathrm{neg}}^{(m)}-\Delta t\nabla_{x}V_{\mathrm{sg}(\theta)}(x_{\mathrm{neg}}^{(m)})+\sqrt{2\Delta t\varepsilon^{(m)}}\eta$triangleright Langevin dynamics given $\varepsilon(t)$
9:         if $x_{\mathrm{neg}}^{(0)}$ initialized from data then
10:             Use $\varepsilon_{\max}$ instead of $\varepsilon^{(m)}$triangleright Use maximum temperature if initialized at the data manifold
11:         end if
12:    end for
13:    $\mathcal{L}_{\mathrm{neg}}\leftarrow\mathbb{E}[V_{\theta}(x_{\mathrm{neg}}^{(M)})]$
14:    $\mathcal{L}_{\mathrm{pos}}\leftarrow\mathbb{E}_{\mathrm{data}}[V_{\theta}(x_{\mathrm{data}})]$$\triangleright$ Compute expected potential energy on the data
15:    $\operatorname{\mathcal{L}_{\mathrm{CD}}}\leftarrow\mathcal{L}_{\mathrm{pos}}-\mathcal{L}_{\mathrm{neg}}$
16:    $\mathcal{L}(\theta)\leftarrow\operatorname{\mathcal{L}_{\mathrm{OT}}}+\lambda_{\mathrm{CD}}\operatorname{\mathcal{L}_{\mathrm{CD}}}$
17:    Update $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta)$triangleright Gradient descent step
18:end for
19:return $\theta$triangleright Trained $\theta$
3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix D.
3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion with one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table 1 and Table 2. Our method surpasses state-of-the-art EBMs, achieving a $50\%$ reduction in FID.
[…]
. Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
[…]
Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories
Ensembles: Diffusion + (one or many) EBMs Diffusion Models
Cooperative DRL (40M params) (Zhu et al., 2024) 9.35 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 8.42
Energy-based Models Flow-based Models
ImprovedCD (Du et al., 2021) 32.48 Flow-matching (Lipman et al., 2023) (196M params) 5.02
CLEL-base (Lee et al., 2023) (7M params) 22.16
CLEL-large (Lee et al., 2023) (32M params) 15.47
Energy Matching (50M params, Ours) 7.67
3.2 Inverse problems
In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w\sim\mathcal{N}\!\bigl{(}0,\sqrt{2}\,\zeta I\bigr{)}$. In this setting, the posterior distribution of $x$ given $y$ is
$p(x|y)\propto\underbrace{\exp\left(-\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update. We detail the algorithm for generating solutions to inverse problems in Algorithm 3 of Appendix B. We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark.
Controlled inpainting.
Suppose we want to recover two images from a masked image while encouraging diverse reconstructions. EBMs allow this by introducing an additional interaction energy, $W(x_{1},x_{2})=-\frac{\|B(x_{1}-x_{2})\|^{2}}{\sigma^{2}}$, where $B$ has ones in the region of interest (focusing diversity there) and zeros elsewhere, and $\sigma$ is a hyper-parameter controlling the interaction’s strength. Specifically, we define $p(x_{1},x_{2}\mid y)\propto p(x_{1}\mid y)\,p(x_{2}\mid y)\,\exp\bigl(-W(x_{1},x_{2})\bigr{missing})$, which gives high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$. This encourages exploring the edges of the posterior rather than just its modes, and with suitable $W$, samples shift toward rare events without needing many draws. To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task. As shown in Figure 2, we start from a partially observed (masked) face and aim to reconstruct two distinct high-fidelity completions.
[…]
Figure 2: Controlled inpainting for diverse reconstructions. On the left is the masked face. On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix $A$ is the dotted blue square (zeros inside, ones outside). By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
Protein inverse design.
In Figure 3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021). Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition. Beyond achieving high fitness, practical inverse design requires generating diverse candidate sequences to ensure robustness in the downstream experimental validation (Jain et al., 2022). We evaluate on two benchmark splits (medium and hard), which correspond to subsets of the original AAV dataset differing in baseline fitness distributions and required mutational distance from known high-performing variants (Kirjner et al., 2024). Leveraging the latent-space representation of VLGPO (Bogensperger et al., 2025), we employ our Energy-Matching Langevin sampler with an inference-time tunable repulsion term, allowing explicit control over the diversity of the designed proteins. This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity. See Section B.2 for experimental details and dataset descriptions.
[…]
Figure 3: Fitness–diversity trade-off for protein inverse design on the AAV Medium (left) and Hard (right) benchmarks. We compare our Energy Matching method (blue), with diversity explicitly controlled by a repulsion strength parameter ($\propto\frac{1}{\sigma^{2}}$), against leading flow-based (purple), score-based (orange), and other non-likelihood methods (black). Fitness measures how well generated sequences satisfy the target property (predicted viral packaging efficiency), while diversity quantifies the average Levenshtein distance between sequences in each generated batch.
3.3 Local intrinsic dimension estimation
Real-world datasets, despite displaying a high number of variables, can often be represented by lower-dimensional manifolds—a concept referred to as the manifold hypothesis (Fefferman et al., 2016). The dimension of such a manifold is called the intrinsic dimension. Estimating the LID at a given point reveals its effective degrees of freedom or directions of variation, offering insight into data complexity and adversarial vulnerabilities. We defer the precise definition to Appendix C.
Diffusion-based approaches.
Recent work leverages pre-trained diffusion models to estimate the LID (Kamkari et al., 2024; Stanczuk et al., 2024) by examining the learned score function. However, because these models do not learn the score at $t=0$ (the data manifold), their estimates become unreliable there. As a result, current methods rely on approximations, for instance, by evaluating the score at small but nonzero times $t=t_{0}$, where the computations remain sufficiently reliable.
Hessian-based LID Estimation.
Unlike diffusion models, EBMs explicitly parametrize the relative data likelihood. This explicit parametrization enables efficient analysis of the curvature of the underlying data manifold – in this example, estimating the LID. To this end, we compute the Hessian matrix $\nabla_{x}^{2}V(x_{\mathrm{data}})$ at a given data point and perform its spectral decomposition. We define near-zero eigenvalues as those whose absolute values lie within a small threshold $\tau$ (in our experiments, we set $\tau=10$ for MNIST (Deng, 2012) and $\tau=2$ for CIFAR-10). The count of near-zero eigenvalues reflects the number of flat directions and thus reveals the local dimension. As shown in Table 3, the LID estimates we obtain exhibit stronger correlations with PNG compression size [^2] PNG is a lossless compression scheme specialized for images and can provide useful guidance when no LID ground truth is available Kamkari et al. (2024). (evaluated on 4096 images) using Spearman’s correlation. Figure 4 offers qualitative illustrations. Our EBM-based approach compares favorably to diffusion-based methods, as it relies on fewer approximations by performing computations exactly on the data manifold rather than merely in its vicinity.
Spearman’s correlation $\uparrow$ MNIST CIFAR-10
ESS (Johnsson et al., 2014) 0.444 0.326
FLIPD (Kamkari et al., 2024) 0.837 0.819
NB (Stanczuk et al., 2024) 0.864 0.894
Energy Matching (Ours) 0.895 0.901
Table 3: Spearman’s correlation coefficients of LID estimates with PNG compression rate. Benchmarks results reported in (Kamkari et al., 2024).
Figure 4: Qualitative results for LID estimation using the Hessian spectrum of $V_{\theta}(x)$. Left: Spectrum for a low-LID image. Right: Spectrum for a high-LID image. The eigenvalues quantify curvature along principal directions (eigenvectors). A degenerate spectrum (many near-zero eigenvalues, marked in red) indicates locally "flat" regions, revealing the LID. Intuitively, higher image complexity often corresponds to a higher LID.
4 Conclusion and limitations
Contributions.
We introduced a generative framework, Energy Matching, that reconciles the advantages of EBMs and OT flow matching models for simulation-free likelihood estimation and efficient high-fidelity generation. Specifically, it:
  • Learns a time-independent scalar potential energy whose gradient drives rapid high-fidelity sampling–surpassing state-of-the-art energy-based models–while also forming a Boltzmann-like density well suitable for controlled generation. All without auxiliary generators.
  • Offers efficient sampling from target data distributions on par with the state-of-the-art, while learning the score at the data manifold with manageable trainable parameters overhead.
  • Offers a simulation-free, principled likelihood estimation framework for solving inverse problems—where additional priors can be easily introduced—and enables the estimation of a data point’s LID with fewer approximations than score-based methods.
Limitations.
First, our method requires the extra gradient computation with respect to input, which can increase GPU memory usage (e.g. 20-40%), in particular during training. Second, for very high-resolution datasets, computing the full Hessian spectrum ($O(N^{3})$) may be impractical; and partial spectrum methods like random projections or iterative solvers may be used instead.
Outlook.
In contrast to the widespread belief, we demonstrated that time-independent curl-free methods for generative flows work and offer an exciting direction for future research. Our Energy Matching approach has the potential to yield novel insights into controlled generation and inverse problems for cancer research Weidner et al. (2024); Balcerak et al. (2024), molecules and proteins Wu et al. (2022); Bilodeau et al. (2022), computational fluid dynamics Gao et al. (2024); Shysheya et al. (2024); Molinaro et al. (2024) and other fields where precise control over generated samples and effective integration of priors or constraints are crucial.
Appendix A Additional details on Energy Matching
In this section, we provide additional studies and visualizations on our method.
A.1 Energy landscape during training
In Figure 5, we visualize how the potential $V_{\theta}(x)$ transitions from a flow-like regime, where the OT loss enforces nearly zero curvature away from the data manifold (a), to an EBM-like regime, where the curvature around the new data geometry (here, two moons) is adaptively increased to approximate $\log p_{\text{data}}(x)$ (b). This two-stage design yields a well-shaped landscape that is both efficient to sample (thanks to a mostly flat potential between clusters) and accurate for density estimation near the data modes. For comparison, (c) shows an EBM trained solely with contrastive divergence, exhibiting sharper but less globally consistent basins.
a) Phase 1 (warm-up):
$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
b) Phase 2 (main training):
$\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}}$
c) Reference EBM:
$\operatorname{\mathcal{L}_{\mathrm{CD}}}$
Figure 5: Visualization of the energy $V_{\theta}(x)$ landscapes driving the samples from eight Gaussians to two moons. See Figure 1 for the 2D perspective. (a) The OT flow loss enforces zero curvature in $V_{\theta}(x)$ along the trajectories to the target. (b) Around the 2 Moons, the curvature of $V_{\theta}(x)$ is adjusted to approximate $\log p_{\text{moons}}(x)\propto V_{\theta}(x)$ while remaining close to the pretrained landscape elsewhere. Combining these objectives yields a potential energy landscape that is both efficient for sampling and representative of the underlying target data distribution. (c) An EBM is shown for comparison, trained using contrastive divergence loss. Visible mode collapse that slows down the equilibration. Less regular landscape away from the data as it needs many simulations to explore it.
A.2 Ablation on the sampling time
Here, we provide ablation studies on CIFAR-10 unconditional generation. Specifically, we first pretrain using $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, and then fine-tune with $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, producing a stable Boltzmann distribution from which one can sample. Figure 6 illustrates the FID as a function of sampling time $\operatorname{\tau_{s}}$ for models trained under these different regimes. In the case of pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$, the quality measure drops (FID increases) sharply when sampling at $\operatorname{\tau_{s}}>1$; this occurs because, once the samples move close to the data manifold, there is no Boltzmann-like potential well to keep them from drifting away. Because the fidelity slope near the data manifold is steep with respect to sampling time, methods lacking explicit time-conditioning can easily overshoot or undershoot, significantly impacting fidelity. This behavior might explain why some models degrade in performance when made time-independent, as recently reported by Sun et al. [2025].
In Figure 6 we also report results for different values of the temperature-switching parameter $\operatorname{\tau^{\ast}}$, which influences the sampling along the paths towards the data manifold (see Equation 3).
11112222333300222244446666888810101010Sampling time $\operatorname{\tau_{s}}$ vs sampling time $\operatorname{\tau_{s}}$$\operatorname{\tau^{\ast}}=0.9,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\tau^{\ast}}=1.0,\ (\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$$\operatorname{\mathcal{L}_{\mathrm{OT}}}$
Figure 6: CIFAR-10 unconditional generation FID vs. sampling time $\operatorname{\tau_{s}}$ when sampling from models trained under different scenarios: pure $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ and combined $(\operatorname{\mathcal{L}_{\mathrm{OT}}}+\operatorname{\mathcal{L}_{\mathrm{CD}}})$, with temperature regime switching parameter $\operatorname{\tau^{\ast}}\in\{0.9,1.0\}$. Lower FID indicates better generative quality. All results for Euler-Heun with $\Delta t=0.01$.
Appendix B Posterior sampling
This section provides additional algorithmic details and experimental settings for inverse problem sampling tasks presented in the main text.
B.1 Algorithm description
The detailed sampling procedure for inverse problems, optionally incorporating an interaction energy term to explicitly control the diversity among generated samples, is provided in Algorithm 3. The Langevin dynamics alternates between evaluating a composite potential $U_{\theta}(x)$, composed of a data fidelity term, a learned energy-based prior $V_{\theta}$, and, optionally, an interaction energy $W(x,x^{\prime})$. This interaction term can, for instance, encourage diversity among parallel chains. Samples are initialized independently from the Gaussian prior and iteratively refined through gradient-based updates. Specific hyperparameters, including the discretization step size $\Delta t$, total sampling time $T$, and the interaction energy strength, are detailed in Appendix D.
Algorithm 3 Sampling for inverse problems (with optional interaction energy)
1:for $m=1$ to $M$ do
2:    Initialize $x_{m}^{(0)}\sim\mathcal{N}(0,I)$triangleright Start each chain from Gaussian prior
3:end for
4:$N\leftarrow\lfloor\operatorname{\tau_{s}}/\Delta t\rfloor$triangleright Number of Langevin steps for sampling time $\operatorname{\tau_{s}}$
5:for $n=0$ to $N-1$ do
6:    for $m=1$ to $M$ do $\triangleright$ Data fidelity + prior + interaction energy
7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}-\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
8:         $\eta\sim\mathcal{N}(0,I)$triangleright Gaussian noise for Langevin step
9:         $x_{m}^{(n+1)}\leftarrow x_{m}^{(n)}-\Delta t\,\nabla_{x}U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\sqrt{2\varepsilon^{(n)}\Delta t}\;\eta$triangleright Langevin step
10:    end for
11:end for
12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples
B.2 AAV inverse design protein generation
We consider protein fitness optimization for the adeno-associated virus (AAV) in the medium and hard data regime as proposed by [Kirjner et al., 2024]. The setup for the latent encoding as well as the backbone architecture for the learned prior in the latent space follows [Bogensperger et al., 2025]. Conditional sampling uses classifier guidance via learned predictor networks $g_{\phi}$ or $\tilde{g}_{\phi}$ to guide samples toward high-fitness regions. The CNN-based networks predicting fitness for AAV sequences are directly sourced from [Kirjner et al., 2024] and are trained only on the limited training data for the medium and hard setup, respectively.
For training we follow Algorithm 1 and Algorithm 2. We then sample 128 sequences using Algorithm 3 (same batch size for all baselines), the hyperparameters are detailed in Appendix D. Generated sequences are evaluated for fitness using the learned oracle from Kirjner et al. [2024]. Further, they are assessed for diversity within the generated sequences and novelty relative to the training set [Jain et al., 2022]. As shown in Table 4, our method achieves state-of-the-art fitness while boosting diversity, and incorporating interaction energy in Algorithm 3 further increases diversity with only a minimal impact on fitness.
Table 4: AAV optimization results. For VLGPO (flow-based) and Energy Matching, medium difficulty uses $g_{\phi}$, hard difficulty uses $\tilde{g}_{\phi}$. Metrics (Fitness↑, Diversity↑, Novelty↑). Reported uncertainty of Fitness is expressed as standard deviation.
AAV medium AAV hard
Method Fitness↑ Diversity↑ Novelty↑ Fitness↑ Diversity↑ Novelty↑
Learning Unnormalized Data Likelihood
Energy-based Models
Energy Matching (Ours) 0.59 (0.0) 5.86 5.0 0.61 (0.0) 4.77 6.7
Energy Matching (+repulsion) (Ours) 0.58 (0.0) 6.22 5.0 0.60 (0.0) 5.22 6.6
Learning Transport/Score Along Noised Trajectories
Flow-based Models
VLGPO [Bogensperger et al., 2025] 0.58 (0.0) 5.58 5.0 0.61 (0.0) 4.29 6.2
Diffusion Models
gg-dWJS [Ikram et al., 2024] 0.48 (0.0) 9.48 4.2 0.33 (0.0) 14.3 5.3
Other Methods
LatProtRL [Lee et al., 2024] 0.57 (0.0) 3.00 5.0 0.57 (0.0) 3.00 5.0
GGS [Kirjner et al., 2024] 0.51 (0.0) 4.00 5.4 0.60 (0.0) 4.50 7.0
AdaLead [Sinai et al., 2020] 0.46 (0.0) 8.50 2.8 0.40 (0.0) 8.53 3.4
CbAS [Brookes et al., 2019] 0.43 (0.0) 12.70 7.2 0.36 (0.0) 14.4 8.6
GWG [Grathwohl et al., 2021] 0.43 (0.1) 6.60 7.7 0.33 (0.0) 12.0 12.2
GFN-AL [Jain et al., 2022] 0.20 (0.1) 9.60 19.4 0.10 (0.1) 11.6 19.6
Appendix C Details on LID estimation
Definition.
To start, we need to introduce the concept of local mass, defined as
[…]
where $p(x)$ is the local density and $B(x_{\text{data}},r)$ is a ball of radius $r$ around $x_{\text{data}}$, i.e. $B(x_{\text{data}},r)=\{x\in\mathbb{R}^{d}:\|x-x_{\mathrm{data}}\|\leq r\}$. The LID is then given by:
$\mathrm{LID}(x_{\mathrm{data}})=d-\lim_{r\to 0}\frac{\log\bigl(M(r)\bigr{missing})}{\log(r)}.$
Intuitively, $M(r)$ measures how much probability mass is concentrated in a ball of radius $r$ around $x_{\mathrm{data}}$. As we shrink this ball, the growth rate of $M(r)$ in terms of $r$ reveals the local dimensional structure of the data.
Assumptions.
In the context of contrastive divergence, we assume that data points $x_{\mathrm{data}}$ lie in well-like regions [Hyvärinen, 2006], i.e.:
[…]
Conceptually, $V(x)$ can be thought of as an energy function; points where $\nabla V(x_{\mathrm{data}})=0$ are near local minima of this energy, and the Hessian $\nabla^{2}V(x_{\mathrm{data}})$ provides information about local curvature (see Figure 4 for a qualitative illustration).
Energy-based density.
We define an energy-based density
$p(x)\;\propto\;\exp\!\Bigl{(}-\tfrac{V(x)}{\varepsilon}\Bigr{)},$
where $\varepsilon$ is a temperature parameter. Near a data point $x_{\mathrm{data}}$ satisfying $\nabla_{x}V(x_{\mathrm{data}})=0$, we can approximate $V(x)$ by its second-order Taylor expansion:
$V(x)\approx V(x_{\mathrm{data}})+\frac{1}{2}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}}).$
Consequently, in view of the assumptions above,
$p(x)\propto\exp\left(-\frac{1}{2\varepsilon}(x-x_{\mathrm{data}})^{\top}\nabla_{x}^{2}V(x_{\mathrm{data}})(x-x_{\mathrm{data}})\right).$
Local mass derivation and the rank of the energy Hessian.
Substituting the local quadratic form of $p(x)$ near $x_{\mathrm{data}}$ into the definition of the local mass $M(r)$, we obtain:
[…]
For small $r$, the dominant contribution depends on the rank of the Hessian $\nabla_{x}^{2}V(x_{\mathrm{data}})$. Let $k=\mathrm{rank}(\nabla_{x}^{2}V(x_{\mathrm{data}}))$. Then, as $r\to 0$, one can show that $M(r)=Cr^{k}$, where $C$ does not depend on $r$. We take the logarithm on both sides and divide by $\log(r)$ to get
$\frac{\log(M(r))}{\log(r)}=\frac{\log(C)+k\log(r)}{\log(r)}=k+\frac{\log(C)}{\log(r)},$
and the second term vanishes as $r\to 0$. Hence,
$\mathrm{LID}(x_{\mathrm{data}})=d-k.$
Practical estimation.
In practice, the LID at a data point $x_{\mathrm{data}}$ can be estimated through the following procedure:
  1. 1.
    Train $V(x)$ with Energy Matching.
  2. 2.
    Compute the Hessian $H=\nabla_{x}^{2}V(x_{\mathrm{data}})$.
  3. 3.
    Perform an eigenvalue decomposition on $H$.
Then the estimated local data-manifold dimension corresponds to the number of directions with negligible curvature (smaller magnitude than some $\tau$).
Appendix D Training details
Below, we detail the training configurations for CIFAR-10, ImageNet 32x32, CelebA, MNIST, and AAV. The gradient of the potential, $\nabla_{x}V(x)$, is computed using automatic differentiation via PyTorch’s autograd [Ansel et al., 2024]. We optimize all models with the Adam optimizer Kingma and Ba [2014], and for CIFAR-10, ImageNet 32x32, and CelebA we maintain an exponential moving average (EMA) of the weights.
While we specifically adopt (i) a one-sided trimmed mean of negative sample energies and (ii) clamping of the contrastive loss for stability, any commonly used EBM technique (e.g., persistent contrastive divergence [Tieleman, 2008], replay buffers, multi-scale negative sampling) could be readily employed.
In our approach, we introduce two hyperparameters, $\alpha$ and $\beta$, to control these stabilizing techniques:
[…]
$\beta$ $=\text{clamp threshold for $\mathcal{L}_{\mathrm{CD}}$ (i.e., we clamp $\mathcal{L}_{\mathrm{CD}}$ to be $\geq-\beta$).}$.
CIFAR-10.
The architecture is shown in Figure 7. We use the same UNet from [Tong et al., 2023] (with fixed $t=0.0$, making it effectively time-independent) followed by a small vision transformer (ViT) [Dosovitskiy et al., 2020] to obtain a scalar output. Hyperparameters are: $\tau_{s}=3.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 200k iterations using Algorithm 1 with EMA $0.9999$ and then 10k more with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=8\times 10^{-3},\alpha=0.1,\beta=-0.05$.
ImageNet 32x32.
The architecture is shown in Figure 7 (same as for CIFAR-10). Hyperparameters are: $\tau_{s}=1.5$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$. We train for 600k iterations using Algorithm 1 with EMA $0.9999$ and then 7k more with Algorithm 2 and EMA $0.999$ on 7xA100. The batch size is 128, learning rate is $8\times 10^{-4}$, $\varepsilon_{\max}=0.01$, and $\lambda_{\mathrm{CD}}=2\times 10^{-4}.$ $\alpha=0.2,\beta=-0.01$.
CelebA.
We scale the CIFAR-10 model by 2$\sim$ 2$\times$ 2 ×; see Figure 8. We keep $\tau_{s}=2.0$, $\tau^{*}=0.9$, $\Delta t=0.01$, $M_{\text{Langevin}}=200$, and train for 250k iterations using Algorithm 1 with EMA $0.9999$ then 7k with Algorithm 2 and EMA $0.999$ on 4xA100. The batch size is 32, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$.
MNIST.
We train on a single RTX6000 directly with Algorithm 2 for 300k iterations using $\Delta t=0.05$, $M_{\text{Langevin}}=50$, $\varepsilon_{\max}=0.05$, and $\lambda_{\mathrm{CD}}=0.25$. The architecture consists of a two-stage CNN followed by a two-layer MLP, comprising a total of 2M parameters.
AAV.
We adopt the one-dimensional CNN architecture as used in [Bogensperger et al., 2025], summing the final-layer activations to obtain the potential. We train for 10k iterations using Algorithm 1 and for 1k iterations using Algorithm 2 on a single GeForce RTX 4090. The batch size is 128, learning rate is $1\times 10^{-4}$, $\varepsilon_{\max}=0.1$, and $\lambda_{\mathrm{CD}}=1\times 10^{-4}$. For Algorithm 3 we use $\tau_{s}=1.7$ for AAV medium and $\tau_{s}=1.3$ for AAV hard, $\tau^{*}=0.9$, $\zeta=0.01$ for AAV medium and $\zeta=0.009$ for AAV hard and $\Delta t=0.01$. We set the target fitness to $y=1$ to aim for the maximum fitness in the generated sequences.
Input(3$\times$32$\times$32)UNet(37M params)Transformer(PatchEmbed+ViT, 12M params)ScalarPotential
Hyperparameter Value
Image size 3$\times$32$\times$32
Base channels (UNet) 128
ResBlocks 2
channel_mult [1, 2, 2, 2]
Attention resolution 16
Attention heads (UNet) 4
Head channels (UNet) 64
Dropout 0.1
Transformer (ViT) Head
Patch size 4
Embedding dim 384
Transformer layers 8
Transformer heads 4
Output scale 1000.0
Figure 7: Diagram of our UNet+Transformer EBM for CIFAR-10 and ImageNet 32x32. A UNet (37M params) processes a 3$\times$32$\times$32 image; its output is fed into a Transformer head (PatchEmbed + 8-layer ViT, 12M params) that produces a scalar potential. Here we employ the identical UNet architecture as in [Tong et al., 2023], but with the time parameter fixed at $t=0$ to render the model time-independent.
Input(3$\times$64$\times$64)UNet(83M params)Transformer(PatchEmbed+ViT, 25M params)ScalarPotential
Hyperparameter Value
Image size 3$\times$64$\times$64
Base channels (UNet) 128
ResBlocks 2
channel_mult [1, 2, 3, 4]
Attention resolution 16
Attention heads (UNet) 4
Head channels (UNet) 64
Dropout 0.1
Transformer (ViT) Head
Patch size 4
Embedding dim 512
Transformer layers 8
Transformer heads 8
Output scale 1000.0
Figure 8: Diagram of our UNet+Transformer EBM for CelebA. A UNet (83M params) processes a 3$\times$64$\times$64 image; its output is fed into a Transformer head (PatchEmbed+8-layer ViT, 25M params) that produces a scalar potential.